solve using the 5 step problem solving process
The sum of three consecutive even integers is 228. Find the integers.

Answer by richwmiller(17219)

don't know what your five step program is.
Anything to do the AA's 12 step program?
n+n+2+n+4=228
3n+6=228
3n=222
n=74
n+2=76
n+4=78
another shorter way
228/3=76 is the middle number
74,76,78